You short sold 500 shares of stock at a price of 36 dollars and an inital margin of 75 percent. if the maintenance margin is 25 percent, at what share price will you receive a margin call? what is your account equity at this stock price?

Margin call price?
Account equity?
Answer #1

Proceeds from short sale is = No of shares short sold * price per share = (500*36) = $ 18,000/.

Initial margin to be maintained is = (18000*75%) = $ 13,500/.

Therefore account value is = 18000+13500 = $ 31,500/.

Margin call price is = ( Account value/(shares sold + maintenance margin shares )

That is = (31500/(500+(500*25%))) = (31500/625) = $ 50.40/.

Margin call price is $ 50.40/.

Account equity is = Account value - margin value

That is = 31500-(500*50.40)

Account equity is = (31500-25200) = $ 6,300/.

Account equity is $ 6,300/.